# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@
[ Scale = MatchLowercase, Path = fonts/, Extension = .ttf , UprightFont = *-Regular, ] 11institutetext: Institute of Theoretical Astrophysics, University of Oslo, P.O.Box 1029 Blindern, N-0315 Oslo, Norway
11email: herman.sletmoen@astro.uio.no
SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver
Herman Sletmoen
@@
SymBoltz is a new Julia package that solves the linear Einstein-Boltzmann equations. It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes and is compatible with automatic differentiation. Cosmological models are built from replaceable physical components in a way that scales well in model space, or alternatively written as one compact system of equations. The modeler should simply write down their equations, and SymBoltz solves them and eliminates friction in the modeling process. Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the analytical Jacobian matrix and its sparsity pattern, and interpolating arbitrary variables from the solution. Implicit solvers integrate the full stiff equations at all times without approximations, which greatly simplifies the code. Performance remains as good as in existing approximation-based codes due to high-order implicit methods that take long time steps, fast generated code, optimal handling of the Jacobian and efficient sparse matrix methods. Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo methods in large parameter spaces, training of emulators, Fisher forecasting and sensitivity analysis. The main features form a synergy that reinforces the design of the code. Results agree with established codes to $0.1\%$ with standard precision. More work is needed to implement additional features and for fast reverse-mode automatic differentiation of scalar loss functions. SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes all contributions.
@@
Cosmology: theory - Methods: numerical
@@
Cosmology is at a crossroads (Freedman, 2017). Despite the enormous success of the $\mathrm{\Lambda CDM}$ model in explaining many observations, the increasing precision of modern observations are revealing tensions between theory and observations. These suggest that the current standard model is not the ultimate truth.
@@
On the theoretical side, this drives a search for the true cosmological model through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016). Efficiently exploring the space of models benefits from numerical tools that are easy to modify. Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes with modular components that are easy to replace. Some of the most important such tools in the cosmological modeling toolbox are linear Einstein-Boltzmann solvers ( ?Boltzmann codes?) like CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a).
@@
On the observational frontier, next-generation surveys like the Square Kilometer Array (Dewdney et al., 2009), Vera C. Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise more precise data. Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that coexist in large $O(100)$-dimensional spaces (Piras et al., 2024). In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman and Gelman, 2011) outperform the traditional Metropolis-Hastings algorithm (Hastings, 1970). They explore parameter space more efficiently, but need both the likelihood and its gradient with respect to parameters. Emulators have become popular to get this from existing non-differentiable codes by training differentiable neural networks to reproduce their output (e.g Bolliet et al., 2024; Bonici et al., 2024). Differentiability is also used in forward-modeling field-level inference with simulations and initial conditions from Boltzmann solvers, for example (Seljak et al., 2017). Automatic differentiation can compute derivatives more accurately and faster than approximate and brute-force finite differences. A differentiable Boltzmann solver is a cornerstone in the modern cosmological modeling toolbox.
@@
SymBoltz is a new Boltzmann solver that aims to meet these needs, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability. At its core, a Boltzmann code solves the Einstein equations for a gravitational theory coupled to some particle species described by Boltzmann equations up to first perturbative order around a homogeneous and isotropic universe. For example, it predicts cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations, or generates initial conditions for non-linear $N$-body simulations of large-scale structure.
@@
This article is structured as follows. Section 1 motivates SymBoltz by reviewing the historical development of Boltzmann codes and methods for computing derivatives. Section 2 describes the structure and main features of SymBoltz and how it differs from other codes. Section 3 shows example usage. Section 4 compares performance to an existing code. Section 5 describes possibilities for future work. Section 6 concludes with the current state of the code. Appendices A, B and C list equations, implementation details, comparison settings and testing methods.
1 History and motivation
@@
Peebles and Yu (1970) were first to numerically integrate a comprehensive set of linear Einstein-Boltzmann equations. Their work was refined over several years, and Ma and Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS. Seljak and Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST. Their line-of-sight integration method has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations. This Fortran codebase has since evolved into CAMB [^1] https://camb.info/ written by Lewis et al. (2000), which is one of the two most used and maintained Boltzmann solvers today. Doran (2005a) ported CMBFAST to C++ with the fork CMBEASY, which structured the code in an object-oriented fashion and improved user-friendliness, but this project is abandoned today.
@@
Lesgourgues (2011a); Blas et al. (2011) started the second major family of Boltzmann solvers with the birth of CLASS [^2] http://class-code.net in C. It improved performance, user-friendliness, code flexibility, ease of modification and control over precision parameters. It was the first cross-check of CAMB from an independent branch and boosted the scientific accuracy of the Planck mission.
@@
Since then a healthy arms race have fueled refinements to CAMB and CLASS. Both codes have spawned many forks for studying alternative models and performing custom calculations. Today they are very well-made, efficient and reliable tools.
@@
Recently, several alternative solvers have appeared on the market. They target new numerical techniques, lack of approximations, differentiability, GPU parallelization, interactivity and symbolic computation. PyCosmo [^3] https://pypi.org/project/PyCosmo/ by Refregier et al. (2018) in Python was the first symbolic-numeric Boltzmann code that generates fast C++ code from user-provided symbolic equations, optimizes the sparse and analytical Jacobian matrix and avoids approximation schemes. Bolt.jl [^4] https://github.com/xzackli/Bolt.jl by Li et al. (2023) in Julia is also approximation-free, supports forward-mode automatic differentiation and uses similar differential equations solvers as SymBoltz. DISCO-EB [^5] https://github.com/ohahn/DISCO-EB by Hahn et al. (2024) in the JAX framework in Python is also differentiable and relaxes approximation schemes to avoid overhead from switching equations on GPUs. SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by all these developments.
1.1 Structure of traditional Boltzmann solvers
The full Einstein-Boltzmann equations are partial differential equations that linearize to ordinary differential equations.
@@ 1.1 Structure of traditional Boltzmann solvers
In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\boldsymbol{u}}/{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p})$ for some initial conditions $\boldsymbol{u}(\tau_{0})$ and parameters $\boldsymbol{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
@@
In practice, however, several properties complicate this task. First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background and perturbations stages. It is common to solve each stage with interpolated input from the previous stage, and joining these can be cumbersome and fragment code. Second, the set of equations is very long and convoluted. Ideally, parts of the equations should be easily replaceable to accommodate different submodels for gravity and particle species. Organizing a code with a suitable modular structure that scales well in model space is hard. Third, wildly different timescales coexist in the equations and make them extremely stiff and intractable to solve with standard explicit ODE solvers. This stiffness must be massaged away in the equations or dealt with numerically in other ways. Fourth, the size of differential equations is very large and increases for more accurate description of relativistic species. Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations. Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and precision must be made.
@@
To overcome these challenges, most Boltzmann solvers are written in low-level high-performance languages like C, C++ and Fortran. They are tightly adapted to the pipeline-like computational structure of the problem (e.g. input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in Lesgourgues (2011a)). This makes sense for programmers, but does not necessarily provide the simplest interface for modelers.
@@
Traditional codes like CAMB and CLASS have nonexistent or thin abstraction layers. To modify them, users must work directly in the low-level numerical code and have a good understanding of its internal structure. For example, to implement a new species, one must often modify the code in many places: input handling for new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining each of these stages, output handling, and so on. This leads to fragmented code where changes related to one species are scattered throughout the code.
@@
This structure scales poorly in model space. As more species or gravitational theories are added, each module becomes intertwined with code from other physical components. Even if deactivated by if statements at runtime, their mere presence in the source code increases its complexity and reduces its readability.
@@
One can alleviate this problem to some extent by instead forking the code for modified models, so the main code base is not polluted. But this just moves the problem. Forking duplicates the entire code base even though only small parts are modified. Forks are often abandoned and do not receive upstream improvements. They are also incompatible with each other unless one merges them into one code, but then one is back to the first problem.
@@
The two-language problem amplifies the problem, as data analysis often happens in slow high-level languages like Python. This shapes Boltzmann solvers to rigid pipelines that must compute everything at once and avoid interception at all costs, in order to maximize performance in the low-level language before passing output back to the high-level language. Some features are really just post-processing of the ODE solutions, but are appended to the pipeline even if peripheral to the core task of an Einstein-Boltzmann solver – solving the Einstein-Boltzmann equations. Features like non-linear boosting and CMB lensing could be made as smaller and interoperating modular packages, but instead become part of ?all-in-one? Boltzmann codes.
@@
These patterns lead to big monolithic Boltzmann solvers. They become increasingly complex as they incorporate more models and features beyond their original core scope. This complexity has even driven development of specialized AI assistants for CLASS (Casas et al., 2025). While existing Boltzmann solvers are impressively well-engineered and such tools are only helpful, we think these are symptoms of unnecessary complexity.
1.2 Boltzmann solver approximation schemes
@@
The Einstein-Boltzmann equations are infamously stiff (Nadkarni-Ghosh and Refregier, 2017). This property of differential equations means their numerical solution is unstable with standard explicit integrators and requires tiny step sizes, making them hard to solve. Stiffness can arise when multiple and very different time scales appear in the same problem. This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid, for example. Stiff equations are practically impossible to integrate with explicit solvers and require special treatment.
For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^7] Here ?approximations? refers to schemes that switch between different equations at different times. It excludes techniques like multipole truncation and line-of-sight integration, which SymBoltz also uses. in the equations (e.g. Doran (2005b); Blas et al. (2011); Lewis (2025)):
  • tight-coupling approximation (TCA),
  • ultra-relativistic fluid approximation (UFA),
  • radiation streaming approximation (RSA),
  • non-cold dark matter fluid approximation (NCDMFA),
@@
They involve switching from one set of equations to another when some control variable measuring the applicability of the approximation crosses a threshold. This can change the variables in the ODE and require reinitializing it. Approximations can help both speed and stability and allow the use of explicit solvers.
@@
However, approximations put much more strain on modelers to derive and validate several versions of the equations in different regimes. This process may have to be repeated when the model is modified. The numerics also become more complicated: timeseries from each ODE solution must be stitched together, each separate ODE system can use different tolerances, ODE integrators must be reinitialized, and so on.
@@
Another way to integrate stiff equations is to use appropriate implicit solvers. PyCosmo first solved the full stiff Einstein-Boltzmann equations with an implicit integrator, followed by Bolt.jl and DISCO-EB. CLASS also has an implicit solver, but does not permit disabling the tight-coupling approximation, for example. Historically, implicit solvers may have been underutilized because they are harder to implement than explicit solvers and have a reputation for being slower, and due to iterative evolution from initial Boltzmann codes that were centered around approximations. However, new life has recently been breathed into the field of implicit solvers, with development of new methods combined with techniques like automatic and symbolic differentiation that make them more feasible and powerful (e.g. Steinebach, 2023; Ekanathan et al., 2025).
1.3 Differentiation methods
@@
Derivatives are important in computing. Cosmological applications are no exception. For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood increases. In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of parameters by minimizing a scalar loss. Some cosmologies are parametrized as boundary-value problems with the shooting method, and use nonlinear root solvers like Newton’s method that need Jacobians. Implicit ODE solvers also use Jacobians to solve for values at the next time step. Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with some uncertainty. Boltzmann solvers save time by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with derivatives with respect to $k$ and $l$. These are all examples where (applications of) Boltzmann solvers need derivatives.
There are at least four ways to compute derivatives.
Manual differentiation is human application of differentiation rules, but is limited to simple expressions and by human error.
@@
Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow like conditional statements and loops that depend on numerical values.
@@
Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times. This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ evaluations ($2n$ using central differences) to compute the gradient of an $n$-variate scalar $f$.
Automatic differentiation can be understood by viewing any computer program as a (big) composite function

### v3

# SymBoltz.jl: A symbolic-numeric, approximation-free, and differentiable linear Einstein–Boltzmann solver

@@
11institutetext: Institute of Theoretical Astrophysics, University of Oslo, PO Box 1029 Blindern, 0315 Oslo, Norway
SymBoltz.jl: A symbolic-numeric, approximation-free, and differentiable linear Einstein–Boltzmann solver
Herman Sletmoen Email: herman.sletmoen@astro.uio.no
@@
SymBoltz is a new Julia package for solving the linear Einstein–Boltzmann equations in cosmology. It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes, and is compatible with automatic differentiation. Cosmological models are built from replaceable physical components in a way that scales well to extended models, or alternatively written as one compact system of equations. The modeler provides their equations, and SymBoltz solves them while reducing friction in the modeling process. Symbolic knowledge enables powerful automation of tasks, such as separating computational stages (e.g., background and perturbations), generating analytical and sparse Jacobian matrices, and interpolating arbitrary variables from the solution. Implicit solvers integrate the full stiff equations at all times without approximations, which greatly simplifies the code. Performance remains comparable to existing approximation-based codes due to efficient high-order implicit methods, fast generated code, optimal handling of the Jacobian, and sparse matrix methods. Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo (MCMC) sampling in large parameter spaces, Fisher forecasting, emulator training, and sensitivity analyses. The main features form a synergy that reinforces the design of the code. Output spectra agree with established codes up to 0.1% with standard precision levels. More work is needed to implement additional features and for fast reverse-mode automatic differentiation of scalar loss functions. SymBoltz is publicly available with single-command installation and extensive documentation. We welcome all contributions to the code from the community.
@@
methods: numerical – cosmology: theory
@@
1 Introduction
@@
Cosmology is at a crossroads (Freedman, 2017), as the precision of modern observations continues to reveal tensions with theory. This suggests that the $\Lambda$ cold dark matter ($\mathrm{\Lambda CDM}$) model is incomplete, despite successfully explaining many observations. On the theoretical side, this situation drives a search for more realistic models through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016). This exploration benefits from access to numerical tools that are easy to modify, encouraging us to relax model-dependent approximations, create user-friendly interfaces, and structure codes with modular components. Linear Einstein–Boltzmann solvers ( ?Boltzmann codes?), such as CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a), are essential tools in the cosmological modeling toolbox.
@@
On the observational frontier, next-generation surveys such as the Square Kilometre Array (Dewdney et al., 2009), Vera C. Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simons Observatory (The Simons Observatory Collaboration, 2019), and Euclid (Euclid Collaboration, 2025) promise more precise data. Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that coexist in large $O(100)$-dimensional spaces (Piras et al., 2024). In these high dimensions, modern Markov chain Monte Carlo (MCMC) methods, such as Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman and Gelman, 2014), outperform the traditional Metropolis–Hastings algorithm (Hastings, 1970). They use both the likelihood and its gradient to explore parameter space faster. Emulators have become a popular proxy to get gradients from nondifferentiable codes by training differentiable neural networks to reproduce their output (e.g., Bolliet et al., 2024; Bonici et al., 2024). Differentiability is also used in forward-modeling field-level inference with simulations and initial conditions from Boltzmann solvers (Seljak et al., 2017). Automatic differentiation can compute derivatives more accurately and faster than approximate and brute-force finite differences. This makes directly differentiable Boltzmann solvers a key addition to the modern cosmological modeling toolkit.
@@
SymBoltz.jl [^1] https://github.com/hersle/SymBoltz.jl is a new Boltzmann solver that aims to meet these needs, featuring a convenient symbolic-numeric interface, approximation-freeness, and differentiability. At its core, a Boltzmann code solves the Einstein equations for a gravitational theory coupled to some particle species described by Boltzmann equations up to first perturbative order around a homogeneous and isotropic universe. For example, it predicts cosmic microwave background (CMB), baryon acoustic oscillation (BAO), or supernova (SN) observations, and generates initial conditions for nonlinear $N$-body simulations of large-scale structure.
@@
This article is structured as follows. Section 1 motivates SymBoltz by reviewing the history of Boltzmann codes and gradient methods. Section 2 presents its structure and main features. Section 3 shows example use. Section 4 compares performance to CLASS. Section 5 outlines paths for future work. Section 6 concludes with the current state of the code. Appendices A, B and C list equations and technical details.
@@
Historically, Peebles and Yu (1970) were the first to numerically integrate a comprehensive set of linear Einstein–Boltzmann equations. Their work was refined over several years, and Ma and Bertschinger (1995) laid the groundwork for modern Boltzmann solvers with the COSMICS code. Shortly after, Seljak and Zaldarriaga (1996) integrated the photon multipoles by parts to reduce thousands of coupled differential equations to independent integral solutions. This line-of-sight integration method was first implemented in their CMBFAST code and has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations. This Fortran codebase has since evolved into CAMB [^2] https://camb.info/ written by Lewis et al. (2000), which is one of the two most used and maintained Boltzmann solvers today. Doran (2005a) ported CMBFAST to C++ with the fork CMBEASY, which structured the code in an object-oriented fashion and improved user-friendliness, but this project is abandoned today.
@@
Lesgourgues (2011a) and Blas et al. (2011) started the second major family of Boltzmann solvers with the birth of CLASS [^3] http://class-code.net in C. It improved performance, user-friendliness, code flexibility, ease of modification, and control over precision parameters. It was the first independent cross-check of CAMB and boosted the scientific accuracy of the Planck mission.
@@
Since then a healthy arms race has fueled refinements to CAMB and CLASS. Both codes have spawned many forks for studying alternative models and performing custom calculations. Today they are very well-made, efficient, and reliable tools.
@@
Recently, several alternative solvers have appeared on the market. They target new numerical techniques, lack of approximations, differentiability, GPU parallelization, interactivity, and symbolic computation. PyCosmo [^4] https://pypi.org/project/PyCosmo/ by Refregier et al. (2018) in Python was the first symbolic-numeric Boltzmann code that generates fast C++ code from user-provided symbolic equations, optimizes the sparse and analytical Jacobian matrix, and avoids approximation schemes. Bolt.jl [^5] https://github.com/xzackli/Bolt.jl by Li et al. (2023) in Julia is also approximation-free, supports forward-mode automatic differentiation, and uses similar differential equations solvers as SymBoltz. DISCO-EB [^6] https://github.com/ohahn/DISCO-EB by Hahn et al. (2024) in the JAX framework in Python is also differentiable and relaxes approximation schemes to avoid overhead from switching equations on GPUs. SymBoltz.jl [^7] https://github.com/hersle/SymBoltz.jl is another Boltzmann code that brings together several of these developments.
@@
The core task of an Einstein–Boltzmann solver is to solve the Einstein–Boltzmann equations for some cosmological model. They are partial differential equations that linearize to ordinary differential equations (ODEs) ${\mathrm{d}\boldsymbol{u}}/{\mathrm{d}\tau}=\boldsymbol{f}(\boldsymbol{u},\boldsymbol{p},\tau)$ for some initial conditions $\boldsymbol{u}(\tau_{i})$ and parameters $\boldsymbol{p}$, including several perturbation wavenumbers. Any output from Einstein–Boltzmann codes is derived from the solution of these ODEs, such as the matter and CMB power spectra.
@@
However, several properties of the equations complicate this task. First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background, perturbations, and line-of-sight integration stages. It is common to solve each stage with interpolated input from the previous stage, and joining these can be cumbersome and fragment code. Second, the set of equations is very long and convoluted. Ideally, parts of the equations should be easily replaceable to accommodate different submodels for gravity and particle species. It is hard to organize a code with a suitable modular structure that scales well in model space. Third, wildly different timescales coexist in the equations, making them extremely stiff and intractable to solve with standard explicit ODE solvers. This stiffness must be massaged away in the equations by approximations or dealt with numerically by implicit ODE solvers. Fourth, the number of differential equations is very large, particularly for accurate descriptions of relativistic species. Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations. Fifth, the perturbations must be solved for many different wavenumbers $k$, and trade-offs between performance and precision must be made.
@@
To overcome these challenges, most Boltzmann solvers are written in low-level high-performance languages such as C, C++, and Fortran. They are tightly adapted to the pipeline-like computational structure of the problem (e.g., input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in Lesgourgues 2011a). This makes sense for programmers, but does not necessarily provide the simplest interface for modelers.
@@
Traditional codes such as CAMB and CLASS have nonexistent or thin abstraction layers. To modify them, users must work directly in the low-level numerical code and have a good understanding of its internal structure. For example, to implement a new species, users must often modify the code in many places: input handling for new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining each of these stages, output handling, and so on. This leads to fragmented code where changes related to one species are scattered throughout the code.
@@
This structure scales poorly in model space. As more species and gravitational theories are added, each module is intertwined with code from other physical components. Even if unused components are deactivated by ?if? statements at runtime, their mere presence in the source code increases its complexity, reduces its readability, and makes it harder to add more models.
@@
We can alleviate this problem to some extent by instead forking the code for modified models, so the main code base is not polluted. But this only shifts the problem. Forking duplicates the entire code base, even though only small parts are modified. Forks are often abandoned and do not receive upstream improvements. They are also incompatible with each other unless we merge them into one code, which reintroduces the first problem.
@@
The two-language problem amplifies this issue, as data analysis often happens in slower high-level languages such as Python. This shapes Boltzmann solvers to rigid pipelines that must compute everything at once and avoid interception at all costs, in order to maximize performance in the low-level language before passing output back to the high-level language. Some features really just post-process the ODE solutions, but are appended to the pipeline even if they are peripheral to the core task of an Einstein–Boltzmann solver. For example, features such as nonlinear boosting and CMB lensing could be outsourced to smaller and interoperating modular packages, but they often become part of ?all-in-one? Boltzmann codes instead.
@@
These patterns lead to big monolithic Boltzmann solvers that become increasingly complex as they incorporate more models and features beyond their original core scope. This complexity has even driven development of specialized AI assistants for CLASS (Casas et al., 2025). While existing Boltzmann solvers are impressively well-engineered and such tools are only helpful, we think these are symptoms of unnecessary complexity.
@@
The Einstein–Boltzmann equations are notoriously stiff (Nadkarni-Ghosh and Refregier, 2017). This property of differential equations means their numerical solution is unstable with standard explicit integrators and requires tiny step sizes, making them hard to solve. Stiffness can arise when multiple and very different time scales appear in the same problem. This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid. Stiff equations are practically impossible to integrate with explicit solvers and require special treatment. For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^8] Here, ?approximations? refers to schemes that switch between different equations at different times. It excludes techniques such as multipole truncation and line-of-sight integration, which SymBoltz also uses. in the equations (e.g., Doran, 2005b; Blas et al., 2011):
  • tight-coupling approximation (TCA);
  • ultrarelativistic fluid approximation (UFA);
  • radiation streaming approximation (RSA);
  • noncold dark matter fluid approximation (NCDMFA);
@@
They involve switching from one set of equations to another when some control variable measuring the applicability of the approximation crosses a threshold. This can change the variables in the ODE and require reinitializing the integration. Approximations enable explicit solvers and can improve both the speed and stability of the solution. However, they put much more load on modelers to derive and validate several versions of the equations in different regimes. This process must generally be repeated with modifications to the model, as they can invalidate the approximations or reintroduce stiffness. They also complicate the numerics, as time series from each ODE solution must be stitched together, each separate ODE system can use different tolerances, ODE integrators must be reinitialized, and so on.
@@
Another way to integrate stiff equations is to use appropriate implicit solvers. PyCosmo first solved the full stiff Einstein–Boltzmann equations with an implicit integrator, followed by Bolt and DISCO-EB. CLASS also has an implicit solver, but does not permit disabling the tight-coupling approximation, for example. Historically, implicit solvers may have been underutilized because they are harder to implement than explicit solvers and have a reputation for being slower, and due to iterative evolution from initial Boltzmann codes that were centered around approximations. However, new life has recently been breathed into the field of implicit methods, with the development of new solvers combined with techniques such as automatic and symbolic differentiation that make them more feasible and powerful (e.g., Steinebach, 2023; Ekanathan et al., 2025).
@@
Derivatives are important in scientific computing and in cosmological applications. For example, algorithms that optimize likelihoods and MCMC samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood increases. In machine learning, the same applies when training neural network emulators for cosmological observables by minimizing a scalar loss function of parameters. Some cosmologies are parameterized as boundary-value problems with the shooting method and use nonlinear root solvers such as Newton’s method, which needs Jacobians. Implicit ODE solvers also use Jacobians to solve for values at the next time step. Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with some uncertainty. Boltzmann solvers save time by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, which can be made more precise with derivatives with respect to $k$ and $l$. These are all examples where (applications of) Boltzmann solvers need derivatives. There are at least four ways to compute derivatives.
@@
Manual differentiation is human application of differentiation rules, but it is limited to simple expressions and by human error. Symbolic differentiation automates this process with computer algebra systems, but it is inherently symbolic and cannot differentiate arbitrary programs with control flow, such as conditional statements and loops that depend on numerical values.
@@
Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times. This can differentiate arbitrary programs, but it is approximate and introduces the step size $\epsilon$ as a hyperparameter that must be tuned for accuracy and stability. It is a brute-force approach that requires $O(n)$ evaluations ($2n$ using central differences) to compute the gradient of a function $f$ of $n$ variables. Automatic differentiation (e.g., Griewank and Walther, 2008) can be understood by viewing any computer program as a (big) composite function